OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the graphing method to solve the following system of equations. (3 points) Part 2: Provide the solution to the system. (2 points) Part 3: Explain what the solution represents on the graph. (1 point) 3x – 6y = 12 9x + 2y = –24

OpenStudy (anonymous):

graph first equation, graph second equation, look at where they meet

OpenStudy (anonymous):

Part 1: First, you would rewrite each equation in y=mx+b form. This is done by isolating the y-variable. Once both equations are in y=mx+b form, use the slope (m) and the y-intercept (b) to graph both lines. Part 2: solution: Multiply each term in the first equation by 3 so that both equations start with 9x. Now subtract the two equations (eliminating the x-term). We end up with -20y = 60, so y = -3 (be careful when subtracting negative integers). Substitute y = -3 into one of the equations to solve for x, and we end up with x = -2. Part 3: The solution (-2, -3) represents the point of intersection of the two lines on the graph. It is the only point that is common to both lines.